During a school lunch, the cafeteria sold two-fifths as many white milks as chocolate milks. If there were 35 milks sold, how many chocolate milks were sold?
Answer by stanbon(75887)

During a school lunch, the cafeteria sold two-fifths as many white milks as chocolate milks. If there were 35 milks sold, how many chocolate milks were sold?
-----
w = (2/5)c
w + c = 35
-----------
Substitute for "w" and solve for "c":
(2/5)c + c = 35
------
(7/5)c = 35
(1/5)c = 5
chocolate = 25
----------------
Cheers,
Stan H.
------------